# A Generalized Sylvester-Fermat-Torricelli problem with application in disaster relief operations by UAVs

### v1

@@ 2.2 Problem Formulation
Let $F\subset\mathbb{R}^{n}$ be a nonempty closed convex set. The Minkowski gauge function associated with set $F$, is defined as follows:
$\rho_{F}(x):=\inf\{t\geq 0\ |\ x\in tF\},\ \ x\in\mathbb{R}^{q}.$ (2.5)
@@
The set F is often referred to as the dynamics set. The Minkowski gauge function, which is common in the literature, plays a key role in convex analysis and its applications [22]. However, for our purposes in this paper, we require a more flexible function, as defined below.
@@
Given a nonempty set $\Omega\subset\mathbb{R}^{q}$, we define the set-based Minkowski gauge function associated with the sets $F$ and $\Omega$ as follows:
$\rho_{F}^{\Omega}(x):=\inf\{t\geq 0\ |\ x\in\Omega+tF\},$ (2.6)
@@
where $\Omega+tF:=\{\omega+tf\ |\ \textnormal{for all }\omega\in\Omega\textnormal{ and }f\in F\}$. In the function $\rho_{F}^{\Omega}$, the sets $F$ and $\Omega$ are referred to as the dynamic set and the reference set, respectively. In simple terms, if we view $F$ as a dynamic set representing possible speeds or directions of motion, then $\rho_{F}^{\Omega}(x)$ represents the minimal time required to reach $x$ starting from at least one point $\omega\in\Omega$. Moreover, $\rho_{F}$ can be interpreted in this context by noting that $\rho_{F}(x)=\rho_{F}^{\{0\}}(x)$. The set-based formulation of the Minkowski gauge function has been studied in the literature, where a slightly different notation—denoted by $T_{\Omega}^{F}(x)$ in (8.7)—is commonly used. Intuitively, $T^{F}_{\Omega}(x)$ represents the minimal time required to reach the set $\Omega$ when starting from the point $x$ and moving according to the dynamics defined by the set $F$  [^1] The set of points reachable from the origin in one unit of time.. Since our application is primarily concerned with the minimal time required to reach a point $x$ from the set $\Omega$ under dynamics $F$, we adopt the definition introduced in (2.6) for clarity and consistency. Lemma 8.13 proves that both definitions of the set-based Minkowski gauge function are equivalent.
@@
With these preliminaries in place, we now proceed to formulate the proposed model. Let the index sets $I_{k}\neq\emptyset$ for $k=1,2,\dots,n$ form a partition of the set $I:=\{1,2,\dots,m\}$, such that $\bigcup_{k=1}^{n}I_{k}=I$ and $I_{k}\cap I_{l}=\emptyset$ for all $k\neq l$ with $k,l\in\{1,2,\dots,n\}$. For $i\in I_{k}$ with $k=1,2,\dots n$, let $F_{i}\subset\mathbb{R}^{q}$ be a compact, convex set containing the origin in its interior point (i.e., $0\in\operatorname{int}F_{i}$). Also, let $\Omega_{i}$ and the constraint set $\Omega_{0}$ be nonempty, closed, convex subsets of $\mathbb{R}^{q}$. Then, the generalized Sylvester-Fermat-Torricelli (SFT) problem is proposed:
$\min S(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (2.7)
where
$S(x):=\max\{\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ k=1,2,...,n\}.$ (2.8)
In the case where $I_{k}=\{k\}$, the generalized SFT problem reduces to a generalized Sylvester problem, which can be stated as follows:
$\min Y(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (2.9)
where
$Y(x):=\max\{\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ i=1,2,3,\dots,n\}.$ (2.10)
Furthermore, by considering $k=1$ and $|I_{1}|=m$, the generalized SFT problem simplifies to a generalized Fermat–Torricelli problem, which is given by:
$\min T(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (2.11)
where
@@
In this paper, we primarily focus on the generalized SFT problem, which encompasses both the generalized Sylvester and Fermat–Torricelli problems as special cases. A detailed discussion of these problems is provided separately in the Appendix. It is important to emphasize that, unlike the generalized Sylvester problem in the literature—which typically assumes a single identical dynamic set in the space ($F_{i}=F$ for $i=1,2,\dots,n$)—the generalized SFT problem accommodates multiple dynamic sets. Moreover, another key generalization is the use of $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)$ instead of a single $\rho_{F_{i}}^{\Omega_{i}}(x)$, which enables modeling of more complex movement patterns, such as round-trip UAV flights under wind conditions between two points.
@@ 3 Properties of the Set-Based Minkowski Gauge Function
In this section, we first review several foundational definitions and then develop key properties of the set-based Minkowski gauge function which are:
@@ 2. $\bullet$
    Convexity of $\rho_{F}^{\Omega}$, which enables the application of tools from convex analysis.
@@ 3. $\bullet$
    Lipschitz continuity of $\rho_{F}^{\Omega}$ which implies that its subdifferential is bounded.
@@ 4. $\bullet$
    Characterization of the subdifferential of $\rho_{F}^{\Omega}$ at both interior points ($x\in\Omega$) and exterior points ($x\notin\Omega$).
@@ 5. $\bullet$
    An upper estimate for the subdifferential of $\rho_{F}^{\Omega}$, offering a useful means for computing its subdifferential at differentiable points.
In the remainder of the section, we introduce the maximal set-based Minkowski gauge function $\bar{\rho}_{F}^{\Omega}$, as defined in (3.6), and establish its fundamental properties.
After presenting the following definitions, we introduce the first property of $\rho_{F}^{\Omega}$. Let $F\subset\mathbb{R}^{q}$ be a nonempty set; the support function associated with $F$ is defined as follows:
$\sigma_{F}(v):=\sup\{\langle v,f\rangle\ |\ f\in F\},\ v\in\mathbb{R}^{q}.$ (3.1)
The $r$-enlargement of a set $\Omega\subset\mathbb{R}^{q}$ is defined as:
$\Omega_{r}:=\{x\in\mathbb{R}^{q}\mid\rho_{F}^{\Omega}(x)\leq r\},\ \ \text{for any }\ r>0.$ (3.2)
As the first property of $\rho_{F}^{\Omega}$, we establish its relationship with the Minkowski gauge function.
Proposition 3.1 (The relationship between $\rho_{F}^{\Omega}$ and $\rho_{F}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed convex set. Then, for a nonempty set $\Omega\subset\mathbb{R}^{q}$, we have:
@@
First, consider the case where $\rho_{F}^{\Omega}(x)<\infty$. Let $\epsilon>0$, and suppose $\rho_{F}^{\Omega}(x)=p$. Choose $t$ such that $p\leq t<p+\epsilon$ and $x\in\Omega+tF$. Then, there exists $\omega\in\Omega$ such that $x-\omega\in tF$, which implies $\rho_{F}(x-\omega)\leq t<p+\epsilon$. Since $\rho_{F}^{\Omega}(x)=p$, it follows that
$\inf_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-\omega)<\rho_{F}^{\Omega}(x)+\epsilon.$
Taking the limit as $\epsilon\to 0$, we obtain
@@
For the reverse inequality, suppose $\inf_{\omega\in\Omega}\rho_{F}(x-\omega)=q$. Let $\alpha>0$, and choose $d$ and $\omega$ such that $q\leq d<q+\alpha$ as well as $x-\omega\in dF$. Then, $x\in\omega+dF\subset\Omega+dF$. Hence, $\rho_{F}^{\Omega}(x)\leq d<q+\alpha$. Taking the limit as $\alpha\to 0$, we get
@@
Combining both inequalities, we conclude that
$\rho_{F}^{\Omega}(x)=\inf_{\omega\in\Omega}\rho_{F}(x-\omega).$
Finally, consider the case where $\rho_{F}^{\Omega}(x)=\infty$. Then, the set $\{t\geq 0\mid x\in\Omega+tF\}=\emptyset$. This implies that for every $\omega\in\Omega$, the set $\{t\geq 0\mid x-\omega\in tF\}=\emptyset$, so $\rho_{F}(x-\omega)=\infty$. Therefore,
$\inf_{\omega\in\Omega}\rho_{F}(x-\omega)=\infty=\rho_{F}^{\Omega}(x).$
For the next property, we prove the convexity of the set-based Minkowski gauge function under the convexity of the reference set in the following proposition.
Proposition 3.2 (Convexity of $\rho_{F}^{\Omega}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed convex set. If $\Omega\subset\mathbb{R}^{q}$ is nonempty and convex, then $\rho_{F}^{\Omega}(.)$ in (2.6) is a convex function.
@@
Given $\epsilon>0$ and select $t_{1}$ and $t_{2}$ such that $\rho_{F}^{\Omega}(x_{i})\leq t_{i}<\rho_{F}^{\Omega}(x_{i})+\epsilon$ as well as $x_{i}\in\Omega+t_{i}f_{i}$ where $f_{i}\in F$ for $i=1,2$. Since $x_{i}-t_{i}f_{i}\in\Omega$ (for $i=1,2$) and convexity of $\Omega$, we have $\lambda(x_{1}-t_{1}f_{1})+(1-\lambda)(x_{2}-t_{2}f_{2})\in\Omega$, for $\lambda\in(0,1)$.
Hence,
@@
By convexity of $F$, $\lambda t_{1}F+(1-\lambda)t_{2}F=(\lambda t_{1}+(1-\lambda)t_{2})F$, so we have
$\lambda x_{1}+(1-\lambda)x_{2}\in\Omega+(\lambda t_{1}+(1-\lambda)t_{2})F\ \Rightarrow\ \rho_{F}^{\Omega}(\lambda x_{1}+(1-\lambda)x_{2})\leq\lambda t_{1}+(1-\lambda)t_{2}.$
By arbitrary choice of $\epsilon$, we have
$\lambda t_{1}+(1-\lambda)t_{2}<\lambda\rho_{F}^{\Omega}(x_{1})+(1-\lambda)\rho_{F}^{\Omega}(x_{2})+\epsilon\Rightarrow\rho_{F}^{\Omega}(\lambda x_{1}+(1-\lambda)x_{2})\leq\lambda\rho_{F}^{\Omega}(x_{1})+(1-\lambda)\rho_{F}^{\Omega}(x_{2}).$
The following proposition establishes the Lipschitz continuity of the function $\rho_{F}^{\Omega}$—a key property for our analysis.
Proposition 3.3 (Lipschitz Continuity of $\rho_{F}^{\Omega}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty, closed, convex set with $0\in$ int $F$. Then for a nonempty $\Omega\subset\mathbb{R}^{q}$, $\rho_{F}^{\Omega}(\cdot)$ is Lipschitz continuous.
@@
The function $\rho_{F}$ is subadditive and positively homogeneous (Theorem 6.14 in [22] provides its proof). Therefore, for any $x,y\in\mathbb{R}^{q}$, we have:
@@
$\Rightarrow\inf_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\inf_{\omega\in\Omega}\rho_{F}(y-\omega)$
$\Rightarrow\inf_{\omega\in\Omega}\rho_{F}(x-\omega)-\inf_{\omega\in\Omega}\rho_{F}(y-\omega)\leq\rho_{F}(x-y)$
$\Rightarrow\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(y)\leq\rho_{F}(x-y).$
Also, for the function $\rho_{F}$, there exists a constant $l\geq 0$ such that $\rho_{F}(x)\leq l\|x\|$ for all $x\in\mathbb{R}^{q}$, as shown in Proposition 6.18 of [22]. Hence, we have:
$\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(y)\leq\rho_{F}(x-y)\leq l\|x-y\|.$
After establishing Lipschitz continuity, we proceed to characterize the subdifferential of the function $\rho_{F}^{\Omega}(x)$ at points $x\in\Omega$ (interior points).
Proposition 3.4 (Subdifferential of $\rho_{F}^{\Omega}$ at interior points).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega$ be a nonempty convex set in $\mathbb{R}^{q}$. Then, the subdifferential of $\rho_{F}^{\Omega}(\cdot)$ at $\bar{x}\in\Omega$ is given by
$\partial\rho_{F}^{\Omega}(\bar{x})=N(\bar{x};\Omega)\cap\mathcal{V}$
where $\mathcal{V}$ is defined as
$\mathcal{V}:=\{v\in\mathbb{R}^{q}\mid\sigma_{F}(v)\leq 1\}.$
@@
Fix any $v\in\partial\rho_{F}^{\Omega}(\bar{x})$ and for $x\in\Omega$, $\rho_{F}^{\Omega}(x)=0$, we get the inequlities,
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})=0.$
Therefore, $v\in N(\bar{x};\Omega)$. For $x\not\in\Omega$, there exists $t>0$ and $f\in F$ such that $\bar{x}+tf=x$ which means $\rho_{F}^{\Omega}(x)\leq t$. Then we have:
$\langle v,\bar{x}+tf-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-0\leq t.$
Therefore, $\sigma_{F}(v)\leq 1$, and $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq N(\bar{x};\Omega)\cap\mathcal{V}$.
To prove the inverse inclusion, fix any $v\in N(\bar{x};\Omega)\cap\mathcal{V}$. Given $\epsilon>0$, for $x\in\mathbb{R}^{q}$, select $t>0$ such that $\rho_{F}^{\Omega}(x)\leq t<\rho_{F}^{\Omega}(x)+\epsilon$. Then, find $\omega\in\Omega$ and $f\in F$ such that $\omega+tf=x$. Therefore,
$\langle v,\omega+tf-\bar{x}\rangle=\langle v,\omega-\bar{x}\rangle+t\langle v,f\rangle\leq t<\rho_{F}^{\Omega}(x)+\epsilon=\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})+\epsilon,$
and $\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})$ as $\epsilon\rightarrow 0$. Hence, $v\in\partial\rho_{F}^{\Omega}(\bar{x})$, and $N(\bar{x};\Omega)\cap\mathcal{V}\subseteq\partial\rho_{F}^{\Omega}(\bar{x}).$
We are now prepared to present the second theorem regarding the subdifferential of the $\rho_{F}^{\Omega}(x)$ function at points $x\notin\Omega$ (exterior points).
Proposition 3.5 (Subdifferential of $\rho_{F}^{\Omega}$ at exterior points).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty convex set. Suppose $\rho_{F}^{\Omega}(\bar{x})=r>0$, where $\bar{x}\notin\Omega$ and $\bar{x}\in\operatorname{dom}\rho_{F}^{\Omega}$. Then the subdifferential of $\rho_{F}^{\Omega}$ at $\bar{x}$ is given by:
$\partial\rho_{F}^{\Omega}(\bar{x})=N(\bar{x};\Omega_{r})\cap\Upsilon\ \ \textnormal{with}\ \ \Upsilon:=\{v\in\mathbb{R}^{q}\ |\ \sigma_{F}(v)=1\}$
@@
Fix $v\in\partial\rho_{F}^{\Omega}(\bar{x})$. We consider three cases:
First, suppose $x\notin\Omega_{r}$. Then, there exist $t>0$ and $f\in F$ such that $\bar{x}+tf=x$, which implies $\rho_{F}^{\Omega}(x)\leq t+r$ by Lemma 8.7. Hence, we obtain the inequality:
@@
which implies $\langle v,f\rangle\leq 1$, and thus $\sigma_{F}(v)\leq 1$.
Second, suppose $x\in\Omega_{r}$. Then the following inequality holds:
@@
which implies $v\in N(\bar{x},\Omega_{r})$.
Third, consider the case $x\in\Omega$. For any $\epsilon>0$, there exist $x\in\Omega$ and $f\in F$ such that $x+df=\bar{x}$, where $r\leq d<r+\epsilon$. Then, we have:
$\langle v,x-\bar{x}\rangle=\langle v,\bar{x}-df-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})$
$\Rightarrow r\leq\langle v,df\rangle<(r+\epsilon)\langle v,f\rangle$
$\Rightarrow\frac{r}{r+\epsilon}<\langle v,f\rangle\leq\sigma_{F}(v).$
As $\epsilon\to 0$, we obtain $\sigma_{F}(v)\geq 1$. Hence, we conclude:

### v2

@@ 2.2 Problem Formulation
Let $F$ be a nonempty, closed, and convex subset of the space $\mathbb{R}^{q}$. The Minkowski gauge function associated with the set $F$, is defined by
$\rho_{F}(x):=\inf\big\{t\geq 0\ \big|\ x\in tF\big\},\ \ x\in\mathbb{R}^{q}.$ (2.4)
@@
The Minkowski gauge (2.4), which is common in the literature, plays a key role in convex analysis and its applications; see, e.g., [mordukhovich2023easy]. However, for our purposes in this paper, we require a more flexible function as defined below.
@@
Given a nonempty set $\Omega\subset\mathbb{R}^{q}$ in addition to $F$, we define the set-based Minkowski gauge function associated with the sets $F$ and $\Omega$ by
$\rho_{F}^{\Omega}(x):=\inf\big\{t\geq 0\ \big|\ x\in\Omega+tF\big\},$ (2.5)
@@
where $\Omega+tF:=\{\omega+tf\ |\ \textnormal{for all }\omega\in\Omega\textnormal{ and }f\in F\}$. In (2.5), the sets $F$ and $\Omega$ are referred to as the (constant) dynamic set and the reference set, respectively. In simple terms, if we view $F$ as a dynamic set representing possible speeds or directions of motion, then $\rho_{F}^{\Omega}(x)$ represents the minimal time required to reach $x$ starting from at least one point $\omega\in\Omega$. Moreover, $\rho_{F}$ can be interpreted in this context by noting that $\rho_{F}(x)=\rho_{F}^{\{0\}}(x)$, see Fig. 1 for a graphical illustration.
Figure 1: (a) illustrates an example of the Minkowski gauge function (2.4) with a circle as its dynamic set. (b) shows an example of the set-based Minkowski gauge function (2.5), where both the reference set and the dynamic set are circles.
@@
With these preliminaries in place, we now proceed to formulate the proposed model. Let the index sets $I_{k}\neq\emptyset$ for $k=1,\dots,n$ form a partition of the set $I:=\{1,\dots,m\}$, such that $\bigcup_{k=1}^{n}I_{k}=I$ and $I_{k}\cap I_{l}=\emptyset$ for all $k\neq l$ with $k,l\in\{1,\dots,n\}$. For $i\in I_{k}$ with $k=1,\dots n$, let $F_{i}\subset\mathbb{R}^{q}$ be a compact and convex set containing the origin in its interior point (i.e., $0\in\operatorname{int}F_{i}$). Suppose also that $\Omega_{i}$ and the constraint set $\Omega_{0}$ are nonempty, closed, and convex in $\mathbb{R}^{q}$. Then the proposed generalized Sylvester-Fermat-Torricelli (SFT) problem is formulated as follows:
$\min S(x)\;\mbox{ subject to }\;x\in\Omega_{0},$ (2.6)
where the cost function in (2.6) is defined by
$S(x):=\max\big\{\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)\ \big|\ k=1,,...,n\big\}.$ (2.7)
In the case where $I_{k}=\{k\}$ and $n=m$, the novel generalized SFT problem is written as
$\min Y(x)\;\mbox{ such that }\;x\in\Omega_{0}$ (2.8)
with the cost function given in the simple maximum form
$Y(x):=\max\big\{\rho_{F_{i}}^{\Omega_{i}}(x)\ \big|\ i=1,\dots,n\big\}.$ (2.9)
This extends the generalized Sylvester problem studied in [mordukhovich2023easy], where only one dynamic set $F$ is under consideration, and where $\rho^{\Omega_{i}}_{F}$ is replaced by the “minimal time function” $T^{F}_{\Omega_{i}}$ that agrees with $\rho^{\Omega_{i}}_{-F}$; see Lemma 8.12 in the Appendix. Furthermore, if $k=1$ and $|I_{1}|=m$ for the cardinality of $I_{1}$, then the proposed generalized SFT problem reduces to the generalized Fermat–Torricelli one stated as follows:
$\min T(x)\;\mbox{ subject to }\;x\in\Omega_{0}$ (2.10)
with the cost function given in the summation form
@@
The latter extends the previous version explored in [mordukhovich2023easy], where only one dynamic set $F$ is present in (2.11), and where $\rho_{F}^{\Omega_{i}}$ is replaced by $T^{F}_{\Omega_{i}}$.
This paper focuses on the generalized SFT problem (2.6), which encompasses both generalized Sylvester and Fermat–Torricelli problems as special cases. Let us emphasize that, unlike the previously considered version of the generalized Sylvester problem with a single identical dynamic set, our new setting in (2.8), (2.9) accommodates multiple dynamic sets important in practical applications. Moreover, another key generalization is using the sum $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)$ instead of a single term $\rho_{F}^{\Omega}(x)$, which enables us to model more complex movement patterns such as, e.g., round-trip UAV flights under wind conditions between two points.
@@ 3 Properties of the Set-Based Minkowski Gauge Function
In this section, we establish some key properties of the set-based Minkowski gauge function that include the following issues:
@@ 2. $\bullet$
    The convexity of $\rho_{F}^{\Omega}$, which enables us to utilize tools of convex analysis.
@@ 3. $\bullet$
    The Lipschitz continuity of $\rho_{F}^{\Omega}$, which implies that its subdifferential is bounded.
@@ 4. $\bullet$
    Characterizing the subdifferential of $\rho_{F}^{\Omega}$ in both settings of in-set points $x\in\Omega$ and out-of-set points $x\notin\Omega$.
@@ 5. $\bullet$
    Deriving an efficient upper estimate for the subdifferential of $\rho_{F}^{\Omega}$.
In what follows, we use the support function
$\sigma_{F}(v):=\sup\big\{\langle v,f\rangle\ \big|\ f\in F\big\},\ v\in\mathbb{R}^{q},$ (3.1)
associated with $F\subset\mathbb{R}^{q}$, and the $r$-enlargement of $\Omega\subset\mathbb{R}^{q}$ defined by
$\Omega_{r}:=\big\{x\in\mathbb{R}^{q}\;\big|\;\rho_{F}^{\Omega}(x)\leq r\big\}\;\text{ for any }\ r>0.$ (3.2)
The first proposition reveals the precise relationship between the set-based Minkowsi gauge function (2.5) and the classical Minkowski gauge (2.4).
Proposition 3.1.
Let $F\subset\mathbb{R}^{q}$ be a nonempty, closed, and convex set. Given an nonempty reference set $\Omega\subset\mathbb{R}^{q}$, we have the relationship
@@
Starting with the case where $\rho_{F}^{\Omega}(x)<\infty$, pick $\varepsilon>0$ and denote $p:=\rho_{F}^{\Omega}(x)$. Choose further $t$ such that $p\leq t<p+\varepsilon$ and $x\in\Omega+tF$. Then there exists $\omega\in\Omega$ with $x-\omega\in tF$, which yields $\rho_{F}(x-\omega)\leq t<p+\varepsilon$. Since $\rho_{F}^{\Omega}(x)=p$, it follows that
$\inf_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-\omega)<\rho_{F}^{\Omega}(x)+\varepsilon.$
Passing to the limit as $\varepsilon\downarrow 0$ gives us the estimate
@@
To verify the reverse inequality, denote $q:=\inf_{\omega\in\Omega}\rho_{F}(x-\omega)$, pick $\alpha>0$, and choose $d$ and $\omega$ such that $q\leq d<q+\alpha$ and that $x-\omega\in dF$. This leads us to $x\in\omega+dF\subset\Omega+dF$, and hence $\rho_{F}^{\Omega}(x)\leq d<q+\alpha$. Passing to the limit as $\alpha\to 0$, we arrive at
@@
Combining both inequalities justifies the claimed representation (3.3) when $\rho_{F}^{\Omega}(x)<\infty$.
It remains to consider the case where $\rho_{F}^{\Omega}(x)=\infty$. Then we get $\{t\geq 0\mid x\in\Omega+tF\}=\emptyset$. This implies that $\{t\geq 0\mid x-\omega\in tF\}=\emptyset$ for every $\omega\in\Omega$ and so $\rho_{F}(x-\omega)=\infty$. Therefore,
$\inf_{\omega\in\Omega}\rho_{F}(x-\omega)=\infty=\rho_{F}^{\Omega}(x),$
which completes the proof of the proposition. ∎
The next proposition addresses the convexity of the set-based Minkowski gauge function.
Proposition 3.2.
Let $F\subset\mathbb{R}^{q}$ be a nonempty, closed, and convex set. If the reference set $\Omega\subset\mathbb{R}^{q}$ is nonempty and convex, then $\rho_{F}^{\Omega}$ in (2.5) is a convex function.
@@
Fix an arbitrary number $\varepsilon>0$ and select $t_{1}$ and $t_{2}$ such that $\rho_{F}^{\Omega}(x_{i})\leq t_{i}<\rho_{F}^{\Omega}(x_{i})+\varepsilon$ and $x_{i}\in\Omega+t_{i}f_{i}$ where $f_{i}\in F$ as $i=1,2$. By $x_{i}-t_{i}f_{i}\in\Omega$ and the convexity of $\Omega$, we have $\lambda(x_{1}-t_{1}f_{1})+(1-\lambda)(x_{2}-t_{2}f_{2})\in\Omega$ for any $\lambda\in(0,1)$. Hence
@@
The convexity of $F$ tells us that $\lambda t_{1}F+(1-\lambda)t_{2}F=(\lambda t_{1}+(1-\lambda)t_{2})F$, and so we have
$\lambda x_{1}+(1-\lambda)x_{2}\in\Omega+(\lambda t_{1}+(1-\lambda)t_{2})F\ \Longrightarrow\ \rho_{F}^{\Omega}(\lambda x_{1}+(1-\lambda)x_{2})\leq\lambda t_{1}+(1-\lambda)t_{2}.$
Since $\varepsilon>0$ was chosen arbitrarily, this yields the implication
$\begin{array}[]{ll}&\lambda t_{1}+(1-\lambda)t_{2}<\lambda\rho_{F}^{\Omega}(x_{1})+(1-\lambda)\rho_{F}^{\Omega}(x_{2})+\varepsilon\\ &\Longrightarrow\rho_{F}^{\Omega}(\lambda x_{1}+(1-\lambda)x_{2})\leq\lambda\rho_{F}^{\Omega}(x_{1})+(1-\lambda)\rho_{F}^{\Omega}(x_{2}),\end{array}$
which readily justifies the claimed convexity of the set-based Minkowski gauge $\rho_{F}^{\Omega}$. ∎
The following assertion establishes the global Lipschitz continuity of the function $\rho_{F}^{\Omega}$ on the entire space $\mathbb{R}^{q}$, which is a key property for our subsequent analysis.
Proposition 3.3.
Let $F\subset\mathbb{R}^{q}$ be a closed and convex set with $0\in{\rm int}\,F$. Then for any nonempty set $\Omega\subset\mathbb{R}^{q}$, we have that $\rho_{F}^{\Omega}(\cdot)$ is Lipschitz continuous on $\mathbb{R}^{q}$.
@@
It is well known that the Minkowski gauge $\rho_{F}$ is subadditive and positively homogeneous; see, e.g., [mordukhovich2023easy, Theorem 6.4]. Therefore, for any $x,y\in\mathbb{R}^{q}$, we get by using Proposition 3.1 that
@@
$\Longrightarrow\inf_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\inf_{\omega\in\Omega}\rho_{F}(y-\omega)$
$\Longrightarrow\inf_{\omega\in\Omega}\rho_{F}(x-\omega)-\inf_{\omega\in\Omega}\rho_{F}(y-\omega)\leq\rho_{F}(x-y)$
$\Longrightarrow\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(y)\leq\rho_{F}(x-y).$
The Lipschitz continuity of $\rho_{F}$ by [mordukhovich2023easy, Proposition 6.18] gives us a constant $l\geq 0$ such that $\rho_{F}(x)\leq l\|x\|$ whenever $x\in\mathbb{R}^{q}$. Therefore,
$\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(y)\leq\rho_{F}(x-y)\leq l\|x-y\|,$
which verifies the global Lipschitz continuity of the set-based Minkowski gauge. ∎
Now we proceed with calculating the subdifferential of the function $\rho_{F}^{\Omega}$ at any point $\bar{x}\in\mathbb{R}^{q}$ starting with the in-set case $\bar{x}\in\Omega$.
Theorem 3.4.
Let $F\subset\mathbb{R}^{q}$ be compact and convex with $0\in\textnormal{int}F$, and let $\Omega$ be nonempty and convex subset of $\mathbb{R}^{q}$. Then the subdifferential of $\rho_{F}^{\Omega}(\cdot)$ at $\bar{x}\in\Omega$ is calculated by
$\partial\rho_{F}^{\Omega}(\bar{x})=N(\bar{x};\Omega)\cap\mathcal{V},$ (3.4)
where $\mathcal{V}$ is defined via the support function (3.1) as
$\mathcal{V}:=\big\{v\in\mathbb{R}^{q}\;\big|\;\sigma_{F}(v)\leq 1\big\}.$
@@
Fix any $v\in\partial\rho_{F}^{\Omega}(\bar{x})$. Then for $x\in\Omega$ with $\rho_{F}^{\Omega}(x)=0$, we get the conditions
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})=0,$
which yield $v\in N(\bar{x};\Omega)$ by the normal cone definition (2.2). Fix $x\not\in\Omega$ and find $t>0$. $f\in F$ with $\bar{x}+tf=x$, which means that $\rho_{F}^{\Omega}(x)\leq t$. Then we have the inequalities
$\langle v,\bar{x}+tf-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-0\leq t$
and thus arrive at $\sigma_{F}(v)\leq 1$ and therefore justify the inclusion $\subset$” in (3.4).
To verify the reverse inclusion in (3.4), fix any $v\in N(\bar{x};\Omega)\cap\mathcal{V}$ and pick an arbitrary number $\varepsilon>0$. Given $x\in\mathbb{R}^{q}$, select $t>0$ such that $\rho_{F}^{\Omega}(x)\leq t<\rho_{F}^{\Omega}(x)+\varepsilon$. Then find $\omega\in\Omega$ and $f\in F$ such that $\omega+tf=x$. Therefore, we get
$\langle v,\omega+tf-\bar{x}\rangle=\langle v,\omega-\bar{x}\rangle+t\langle v,f\rangle\leq t<\rho_{F}^{\Omega}(x)+\varepsilon=\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})+\varepsilon.$
Passing to the limit therein as $\varepsilon\downarrow 0$ gives us $\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})$, which justifies the inclusion $\supset$” in (3.4) and thus completes the proof of the theorem. ∎
The next theorem provides the subdifferential calculation for $\rho_{F}^{\Omega}$ at out-of-set points $\bar{x}\notin\Omega$.
Theorem 3.5.
Let $F\subset\mathbb{R}^{q}$ be a compact and convex set with $0\in\textnormal{int}F$, let $\Omega\subset\mathbb{R}^{q}$ be nonempty and convex, and let $\bar{x}\notin\Omega$. Denote $r:=\rho_{F}^{\Omega}(\bar{x})$ and consider the $r$-enlargement (3.2). Then the subdifferential of $\rho_{F}^{\Omega}$ at $\bar{x}$ is calculated by
$\partial\rho_{F}^{\Omega}(\bar{x})=N(\bar{x};\Omega_{r})\cap\Upsilon\ \ \textnormal{with}\ \ \Upsilon:=\{v\in\mathbb{R}^{q}\ |\ \sigma_{F}(v)=1\}$ (3.5)
@@
Fix $v\in\partial\rho_{F}^{\Omega}(\bar{x})$ and consider the following three cases.
(a) Suppose that $x\notin\Omega_{r}$. Then there exist $t>0$ and $f\in F$ such that $\bar{x}+tf=x$, which yields $\rho_{F}^{\Omega}(x)\leq t+r$ by Lemma 8.6 in the Appendix. This brings us to the conditions
@@
which imply in turn that $\langle v,f\rangle\leq 1$, and thus $\sigma_{F}(v)\leq 1$.
(b) Supposing that $x\in\Omega_{r}$, we get
@@
which tells us that $v\in N(\bar{x},\Omega_{r})$.
(c) Consider finally the case where $x\in\Omega$. Pick any $\varepsilon>0$ and find $x\in\Omega$ and $f\in F$ such that $x+df=\bar{x}$, where $r\leq d<r+\varepsilon$. Then we have the implications
$\begin{array}[]{ll}&\langle v,x-\bar{x}\rangle=\langle v,\bar{x}-df-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})\\ &\Longrightarrow r\leq\langle v,df\rangle<(r+\varepsilon)\langle v,f\rangle\\ &\Longrightarrow\displaystyle\frac{r}{r+\varepsilon}<\langle v,f\rangle\leq\sigma_{F}(v).\end{array}$
[…]